OpenStudy (jiteshmeghwal9):

If \(\sqrt{14+6\sqrt{5}}=a+\sqrt{b}\), then find the value of \(a+b\) :-

OpenStudy (anonymous):

answer is 3+sqrt5

OpenStudy (anonymous):

so a+b=3+5=8

OpenStudy (jiteshmeghwal9):

Will u plz show ur work step by step ?

OpenStudy (anonymous):

\[\sqrt{14+6\sqrt{5}}=\sqrt{14+2\sqrt{45}}\] you must find two number. their multipication must be 45 and their sum must be 14. so they are 9 and 5. sqrt9=3 and sqrt5. ok??

OpenStudy (jiteshmeghwal9):

hm.yaa

OpenStudy (anonymous):

mmm i think you thank me :)

OpenStudy (jiteshmeghwal9):

ohh yeah i just forgot it. thanx & sorry too bro ;)

OpenStudy (anonymous):

welcome bro :)